OpenStudy (anonymous):

Choose the correct description of the end behavior of the function y = 0.5x. A) As x → ∞, y → 0. B) As x → -∞, y → 0. C) As x → ∞, y → ∞. D) As x → -∞, y → -∞. c is wrong

OpenStudy (anonymous):

Are their restrictions?

OpenStudy (anonymous):

there*

OpenStudy (anonymous):

no it doesnt list any but i got it wrong with c so i am a little confused

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}0.5x=\infty\] \[\lim_{x \rightarrow \ - \infty}0.5x=\ - \infty\] Because the graph of the line y=0.5x goes in both directions, in truth, both C) and D) are correct.

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

it was actually A